OpenStudy (anonymous):

Please help. Solve for x in the proportion 3/x+2 = 6/x-1 x = −5 x = 1 x = 4 x = 5

jimthompson5910 (jim_thompson5910):

The equation is \[\large \frac{3}{x+2} = \frac{6}{x-1}\] right?

OpenStudy (anonymous):

yes:)

jimthompson5910 (jim_thompson5910):

ok thank you for the clarification

jimthompson5910 (jim_thompson5910):

I'll get you started \[\large \frac{3}{x+2} = \frac{6}{x-1}\] \[\large 3(x-1) = 6(x+2)\] \[\large 3x-3 = 6x+12\] Can you take it from here?

OpenStudy (anonymous):

i got x=-5

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

you're welcome